# Harm Amplification in Text-to-Image Models

### v1

@@ 6 Discussion & Conclusion
In this paper, we provide a formal definition of harm amplification in T2I models as occurring when the generated output image reflects more harmful representations that were not explicit in the text input (Section 3). Using two specific kinds of harmful representations (sexually explicit content and violent content), we showed how to quantify harm amplification by providing three methods that use safety classifier harm scores and CLIP embeddings (Section 4). We then empirically illustrated the application of these methodologies on a measurements dataset that simulated real-world model deployment and evaluated these methods on an independent human annotated dataset (Section 5).
6.1 Using Data-Driven Methods to Quantify Harm Amplification
Each of our three data-driven methods for quantifying harm amplification offers its own strengths and limitations. The Distribution-Based Thresholds method may be most appropriate if text and image harm scores are not well calibrated. However, with this method, we assume that image harm scores within each bucket follow a normal distribution. If this assumption is violated, then thresholds for each bucket may not be effective in quantifying when harm amplification occurs. The Bucket Flip method is the easiest method to employ, but best performance is achieved if the text and image harm scores are well calibrated – otherwise, a bucket flip may not necessarily mean that harm is amplified. Another limitation for this method is that (text, image) pairs with close harm scores may be positioned in different buckets which might be assessed as harm amplification instances even when the actual difference in scores is small. While the Image-Text Co-embedding method performs well in a low-resource setting, a fine-tuned model using CLIP embeddings specifically trained to detect harm is likely to yield better results. As we observe in the methods using safety classifiers, specific use-case models can dramatically improve performance when detecting instances of harm amplification.
Across all three methods, we assume the harm amplification criterion derived from the measurements dataset will generalize to new data generated by our measurement model (Stable Diffusion 2.1) as well as new data generated by other T2I models. This assumption holds if there is similar harm amplification rates across different T2I models. Although we observe good performance in our evaluation dataset that encompasses data generated from several different models, further research should explore the generalizability of harm amplification methods in cross-domain T2I models.
6.2 Reduced Reliance on Human Annotation
@@
Our novel framework significantly reduces the reliance on human annotations, which may be especially helpful in low-resource settings. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale. Of course, machine annotations — akin to human annotations Davani et al. (2022); Denton et al. (2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
6.3 Gender Disparities in Harm Amplification
@@
In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups. By identifying gender related disparities, we contribute to existing work Luccioni et al. (2023); Naik and Nushi (2023) demonstrating gender stereotypes in T2I models. The durability of gender stereotypes in AI systems Buolamwini and Gebru (2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes. While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
6.4 Future Work
@@
7 Acknowledgements
We thank Kathy Meier-Hellstern, Alexander Vorontsov, and Vinodkumar Prabhakaran for their feedback and suggestions. We also thank Auriel Wright, Candice Schumann, Lora Aroyo, and Alicia Parrish for discussions on analysis and writing.
Appendix A Experimental Methods Applied on the Measurement Dataset
@@
Appendix B Evaluation Results on Nibbler Data
Figure 3: Confusion Matrix for Method 1: Distribution-based thresholds on the Nibbler evaluation dataset. Using the thresholds for this method applied on the measurement dataset, we evaluated these thresholds on the Nibbler data with human annotations serving as ground truth. In the left figure, we see the confusion matrix for harm amplification predictions for sexually explicit content in Nibbler (negative being no harm amplification occurred and positive being that harm amplification occurred). In the right figure, we show the confusion matrix for harm amplification predictions for violent content in Nibbler.
Figure 4: Confusion Matrix for Method 2: Bucket Flip for for Text and Image Score Bucket on the Nibbler evaluation dataset. With this method applied on the measurement dataset, we evaluated the bucket flip criteria on the Nibbler data with human annotations serving as ground truth. The left figure shows the confusion matrix for sexually explicit content and the right figure shows the confusion matrix for violent content
Figure 5: Comparing the PR curves with different amount of terms in Method 3. For sexually explicit content, we show no notable difference between the PR curves for 3 terms ("porn", "sexual", "nudity") and 10 terms ("porn", "sexual", "nudity", "erotic", "sex", "breasts", "penis", "vagina", "erotica", and "indecency"). AUC difference was also minimal (3 terms: 0.846, 10 terms: 0.848). For violent content, we again see minimal difference between the PR curves for 3 terms ("violent", "weapons", "injury") and 10 terms ("violent", "weapons", "injury", "gore", "blood", "death", "gruesome", "pain", "wound", "brutality"). Again, AUC difference between 3 terms (0.412) and 10 terms (0.397) was small.
Figure 6: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right) for the final 3 terms used.
Appendix C Gender Fairness Analysis
Table 3 shows the efficacy of our proposed methods across both the perceived gender expressions for the harm amplification of sexually explicit content. Please note, we focus on sexually explicit harm only as the harm amplification rates were not statistically different for violent content across the two perceived gender expressions.
@@ Appendix C Gender Fairness Analysis
Distribution-Based Thresholds 0.978 0.971 0.974 0.970 0.970 0.970
Bucket Flip 0.985 0.964 0.974 0.753 0.865 0.805
Image-Text Co-embedding 0.978 0.619 0.758 0.760 0.567 0.650

### v2

@@ Discussion
In this paper, we provide the first formalization of harm amplification in T2I models, defining it as occurring when the generated output image reflects more harmful representations that were not explicit in the text input. Using two specific kinds of harmful representations (sexually explicit content and violent content), we provided a framework of different methods to identify instances of harm amplification under different resource and deployment scenarios. We then empirically illustrated the application of this framework on an independent human annotated dataset.
Reduced Reliance on Human Annotation
@@
Our framework significantly reduces the reliance on human annotations. By primarily using machine annotations, we enable scalability making the measurement process adaptable to larger datasets. While human annotations are valuable for evaluating the method’s efficacy on held out datasets, they are not required for the core measurement process. Furthermore, our framework demonstrates its ability to measure harm amplification without the need for explicit adversarial data generation or red teaming, which are often resource intensive and uncomprehensive. Thus, our proposed framework not only offers robust methodologies for measuring harm amplification but also addresses practical challenges that researchers may encounter when deploying models at scale. Of course, machine annotations — akin to human annotations (Davani, Díaz, and Prabhakaran 2022; Denton et al. 2021) — can also reflect social biases. These biases may stem from the subjective nature of defining ground truth in safety. Further research could explore methodologies to more inclusively capture safety in annotations, taking into account diverse perspectives of what safety harms mean.
Gender and Harm Amplification
@@
In analyzing rates of harm amplification in the evaluation dataset, we identified gender-related disparities, particularly for sexually explicit content. Our findings revealed a higher prevalence of oversexualization in perceived female depicting images compared to perceived male depicting images. These results highlight crucial safety concerns regarding the exacerbation of harmful stereotypes relating to gender. Our work emphasizes the importance of both addressing harm amplification in T2I systems and understanding its impacts in different sociodemographic groups. By identifying gender related disparities, we contribute to existing work (Luccioni et al. 2023; Naik and Nushi 2023) demonstrating gender stereotypes in T2I models. The durability of gender stereotypes in AI systems (Buolamwini and Gebru 2018) necessitates greater attention to development of sociotechnical strategies to prevent the perpetuation and amplification of harmful stereotypes. While we focused on gender, other axes of inequality, such as race, ethnicity, disability, sexuality, and culture, among others, should be explored.
Future Work
@@
Appendix A Appendix
@@
Sexually Explicit
porn sexual
nude sex
18+ naked
nsfw dick
vagina explicit content
uncensored fuck
nipples naked person
sexy
Violence
violence gore
blood attack
bloodshed war
horror fight
weapons injury
death pain
wound brutality
harm
Table 2: Harm concepts for Method 3: Image-Text Co-embedding. These harm concepts were meant to approximate the embedding of sexually explicit and violence.
Figure 4: Confusion matrix for each method (left: Distribution-Based Thresholds, middle: Bucket Flip, right: Image-Text Co-embedding) evaluated on the Nibbler dataset.
Figure 5: Precision-recall curves for the Image-Text Co-embedding method for sexually explicit content (left) and violent content (right).
@@
Distribution-Based Thresholds 0.858 0.979 0.915 0.792 0.987 0.878
Bucket Flip 0.913 0.932 0.923 0.860 0.961 0.908
Image-Text Co-embedding 0.656 0.911 0.763 0.308 0.844 0.451